Tommy is seven years older than Michael. In five years, Tommy will be twice as old is Michael. How old are they now?
Answer by jorel555(1290)

Let Michael be m. Then Tommy is m+7. So:
2(m+5)=m+7+5
2m+10=m+12
m=2
Michael is 2; Tommy is 9. ☺☺☺☺